The length of a large rectangular room is to be 3 m more than twice its width. If the perimeter is to be 36 m, find the length and width of the room.

Answer by checkley79(3341)

L=2W+3
P=2L+2W
36=2(2W+3)+2W
36=4W+6+2W
36=6W+6
6W=36-6
6W=30
W=30/6
W=5 M. FOR THE WIDTH
L=2*5+3
L=10+3=13 M. FOR THE LENGTH
PROOF:
36=2*13+2*5
36=26+10
36=36